OpenStudy (anonymous):

Find the area between y=7sinx and y=10cosx over the interval[ 0,3.14] Sketch the curves if necessary. A=

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi}10\cos x-7\sin x\]